# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

### v3

@@ 7 Conclusion
We propose ChOiRe, a four-step solution framework for individual opinion prediction via differentiating the utilization of user’s explicit versus implicit personae. We further introduce Chain-of-opinion reasoning and answer consistency over variable numbers of input implicit personae guiding the models to derive thorough predictions. ChOiRe achieves strong SOTA results with limited inference calls, demonstrating its strong effectiveness. Additionally, Steps (i) and (ii) of ChOiRe significantly improve the fine-tuning of opinion-aligned models. We strongly suggest that our method should only be used for positive moral intents, avoiding making LLMs echo chambers Vicario et al. (2016). In the future, we will focus on developing frameworks that utilize explicit and implicit personae more efficiently.
@@
One limitation of our proposed ChOiRe framework is that it requires the LLMs to have a good capability in following human instructions to solve tasks such as selecting explicit personae, ranking historical opinions, and explaining personae and opinions one by one via CoO. However, we foresee that this limitation is going to be overcome by cutting-edge AI language models, in the present and near future. Additionally, our method also utilizes user’s personal information from explicit and implicit personae, which may be sensitive to some audiences and not be fully available in the real world. However, to what extent is the personal information provided, our ChOiRe is still able to offer reasonable opinion predictions since it is not constrained by the number of provided explicit personae, or the number of user historical opinions.
@@
A personalized LLM allows users to reinforce their existing beliefs and potentially amplify biased or unethical perspectives, leading to the creation of echo chambers Vicario et al. (2016). This can ultimately harm users by reinforcing polarized or undesirable views. To mitigate this issue, the Chain-of-Opinion (CoO) reasoning from our proposed ChOiRe involves presenting user demography or ideology group responses alongside personalized answers. Additionally, CoO can encourage users to reflect on their previous viewpoints.
@@
Through human evaluations, we observe that our proposed method does not generate any discriminatory, insulting responses. We validate the intermediate steps of our proposed ChOiRe by human evaluation which involves manual labor. We hire annotators to score, and the hourly pay is set to \$$\$15$, which is higher than the local statutory minimum wage. Therefore, we do not anticipate any major ethical concerns raising from human evaluations.
Appendix A Baselines
A.1 Baselines Implementation Details
@@
Model ChatGPT/ChatGPT-Inst/GPT-4/Mistral
Guns Auto. & driverless vehicles Views on gender Com. types & sex. harassment Race
W/o persona 53.07/37.30/——-/30.48 47.73/48.26/——-/41.72 50.53/42.94/——-/37.39 47.73/41.67/——-/29.34 41.95/45.28/——-/37.55
DIO-top8 53.87/57.00/60.39/44.73 45.33/44.78/53.22/41.72 53.21/52.15/63.73/40.09 43.47/45.24/42.86/35.45 43.06/44.65/55.17/41.11
DIO-top8 + CoT 54.55/52.33/——-/55.48 47.22/46.77/——-/49.00 48.11/57.67/——-/54.28 42.39/42.26/——-/42.01 45.63/43.40/——-/49.78
DIO-top8 + SC 54.40/52.85/——-/56.57 43.73/48.26/——-/52.31 55.61/56.44/——-/56.30 45.33/40.48/——-/42.01 45.00/43.40/——-/50.00
ChOiRe 57.06/58.21/63.37/58.00 49.25/51.92/50.00/53.75 59.23/53.07/71.43/57.78 39.88/44.14/47.96/42.08 42.77/47.28/50.57/51.44
Gender & Leadership America in 2050 Trust in science Biomedical & food issues Misinformation
W/o persona 53.13/50.83/——-/43.51 39.73/39.13/——-/41.95 50.40/47.29/——-/48.34 53.87/53.63/——-/53.21 46.93/40.38/——-/53.63
DIO-top8 48.27/54.70/65.55/50.23 46.93/46.20/43.70/35.14 54.93/61.58/61.54/51.65 52.27/55.86/58.03/52.78 49.33/52.11/52.71/50.77
DIO-top8 + CoT 48.58/50.83/——-/55.79 43.05/48.91/——-/43.76 54.10/65.02/——-/58.28 56.91/57.54/——-/57.08 49.57/53.99/——-/53.19
DIO-top8 + SC 49.07/53.60/——-/57.87 45.87/47.83/——-/46.03 56.27/65.52/——-/58.94 53.07/57.54/——-/58.58 45.00/53.52/——-/53.85
ChOiRe 52.22/57.78/63.03/57.87 49.46/48.99/45.37/47.50 56.43/55.50/68.46/60.37 54.75/57.26/61.61/58.58 46.45/53.62/57.36/53.85
Privacy & Surveilance Family & Relationships Economic inequality Global attitudes Political views
W/o persona 43.24/40.28/——-/33.64 47.06/44.36/——-/46.08 43.67/49.15/——-/34.07 46.13/46.71/——-/40.42 40.80/48.95/——-/46.20
DIO-top8 53.24/47.22/47.73/43.31 57.22/57.89/62.50/47.42 45.60/51.98/63.81/41.87 49.60/57.23/66.67/41.27 56.80/46.85/62.07/44.13
DIO-top8 + CoT 53.38/47.22/——-/56.91 59.57/55.64/——-/54.36 47.65/51.98/——-/51.45 46.42/56.58/——-/51.06 53.30/45.45/——-/50.80
DIO-top8 + SC 54.05/47.22/——-/58.06 55.35/54.89/——-/57.04 46.13/51.98/——-/52.76 46.42/55.26/——-/52.89 57.33/47.55/——-/51.67
ChOiRe 54.29/53.33/52.27/58.06 60.00/58.77/63.89/58.50 52.33/50.13/64.76/51.89 44.74/55.26/64.58/52.76 51.05/53.74/67.82/53.34
Table 5: Fine-grained accuracy results of ChatGPT/ChatGPT-Instruct/GPT-4/Mistral. DIO stands for Demographic + Ideology + Opinions (section 4).
@@
Appendix B Extra Analysis
B.1 Additional Baseline Comparisons
@@
In this section, we compare ChOiRe’s FEA and LLMtop-$K$ steps with two simple variants outlined in table 6. Given ChatGPT and Mistral’s strong performance with just $4.79/12$ and $8.83/12$ explicit persona attributes, a crucial question arises: Can comparable performance be achieved by randomly selecting $5/12$ and $9/12$ explicit persona attributes instead of relying on LLMs? The first variant, DIO-top8 + Random FEA, involves randomly selecting $5/12$ and $9/12$ explicit persona attributes. The second variant entails randomly selecting $8$ implicit persona opinions instead of using ChOiRe’s LLMtop-$K$ step. From table 6, we find that randomly selecting explicit persona attributes significantly harms the performance of both models due to the removal of important attributes. Additionally, randomly selecting $8$ implicit persona opinions also adversely affects the models, particularly ChatGPT. These observations underscore the effectiveness and importance of ChOiRe’s FEA and LLMtop-$K$ steps.
B.2 Fine-grained Results of ChatGPT, ChatGPT-Instruct, and GPT-4
Table 5 presents the fine-grained results of ChOiRe and baselines for ChatGPT, ChatGPT-Instruct, and GPT-4.
B.3 ChOiRe versus Self-Consistency
Fig. 3 presents the improvements per topic comparision between ChOiRe and the SOTA baseline with Self-Consistency Wang et al. (2023b) DIO-top8 + SC. We observe that ChOiRe improves $11/15$ topics for both ChatGPT and ChatGPT-Instruct.
Figure 3: % of improvements over the SOTA method (DIO-top8 + SC) with ChatGPT-Instruct (left) and ChatGPT (right).
B.4 Accuracy Distribution over Users
Fig. 4 shows the accuracy distribution over users of ChOiRe with ChatGPT. We see that the peak accuracy is at 0.5 for the majority, with a few users scoring zero and over 20 achieving perfection.
Figure 4: Frequency distribution of accuracy over users by ChOiRe.
B.5 Top-3 Removed Explicit Personae Attributes
Table 7 presents the top-3 explicit personae that got removed the most by the LLMs. Among the removed personae, "Citizenship" appears to be the highest-frequency one across models, followed by "Race".
@@ B.5 Top-3 Removed Explicit Personae Attributes
B.6 Ranking Consistency for LLMtop-$K$ Step
We record the average Overlap coefficient Vijaymeena and Kavitha (2016) among $5$ ranking outputs from $5$ input strategies in fig. 5. The performance of those input strategies is further presented in table 8 on $300$ random samples.
@@ B.6 Ranking Consistency for LLMtop-$K$ Step
Model Method Semantic Similarity Order Seed = 2024 Seed = 5 Seed = 2000 Seed = 15 Std
ChatGPT DIO-LLMtop8 - 51.03 50.95 51.11 - 0.0652
Mistral-7B-Instruct-v0.2 DIO-LLMtop8 - 45.86 45.55 45.36 - 0.2060
@@
B.7 Kendall’s Tau Scores for Ranking Agreements
@@
Figure 6: Left: Ranking agreements between ChatGPT top-$K$ and semantic similarity top-$K$. Right: Between ChatGPT-Instruct top-$K$ and semantic similarity top-$K$. One example that has a high disagreement score is shown in section E.2.
Figure 7: Ranking agreements between Mistral top-$K$ and semantic similarity top-$K$.
@@
Fig. 6 shows our ranking agreement scores between ChatGPT and Semantic similarity metric (Left), and ChatGPT-Instruct and Semantic similarity metric (Right). We observe that the two ranking orders have minimal monotonous relations with means approximating 0 and low standard deviations. More specifically, with ChatGPT, the maximum agreement is 0.6000 while the minimum is -0.5895 and the Kurtosis is -0.2173. For ChatGPT-Instruct, the maximum is slightly lower with 0.5473, while the minimum is -0.7368 which is smaller ChatGPT, and the Kurtosis is -0.1017.
B.8 Consistency Scores
Table 9 presents the exact consistency scores for the fig. 2. Besides CoO consistently outperforming CoT, we also observe that when the temperature is increased, the consistency score is decreased, which is intuitive.
Model Temperature Consistency Score (%)
DIO-top8 + CoT 0.3 84
DIO-top8 + CoO 0.3 86
DIO-top8 + CoT 0.6 79
DIO-top8 + CoO 0.6 82
DIO-top8 + CoT 0.9 58
DIO-top8 + CoO 0.9 60
Table 9: Consistency scores of CoT and CoO on $100$ random question-answer pairs. We sample $5$ answers per question and measure the % of questions that have all $5$ identical answers.
Appendix C Prompts and Prompts Analysis
C.1 Prompt Templates for Filtering Explicit Personae
We present the prompt template for selecting relevant explicit personae for answering the question below. The template is hand-crafted and we use Chain-of-Thought (CoT) prompting Kojima et al. (2022) via adding "answer the above question step by step".

### v5

@@ 7 Conclusion
This paper identifies two major challenges in aligning LLMs with human opinions via personae: noisy personae and ineffective reasoning strategies over personae. To address these, we propose COO, a novel four-step framework in the light of the Value-Belief-Norm theory: the first two steps tackle the noise issue using LLMs as analysts, while the last two enhance reasoning through the novel VBN reasoning. COO significantly improves LLM prompting and fine-tuning, demonstrating high generalizability in scenarios with missing personae and potential applications in other personalization tasks.
Acknowledgements
This research project is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-010-SGIL), the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207), and the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2022-005). Do Xuan Long is supported by the A*STAR Computing and Information Science (ACIS) Scholarship.
@@ Acknowledgements
One limitation of COO is that it requires the LLMs to be well capable of following human instructions to solve tasks such as selecting explicit personae, ranking historical opinions, and explaining personae and opinions VBN reasoning. However, we foresee that this limitation is going to be overcome by cutting-edge AI language models, in the present and near future. Additionally, COO utilizes user’s personal information from explicit and implicit personae, which may be sensitive to some audiences and not be fully available in the real world. However, to what extent is the personal information provided, our COO is still able to offer reasonable opinion predictions since it is not constrained by the number of provided explicit personae, or the number of user historical opinions (see Section 6.1). Finally, for fine-tuning, COO currently leverages ChatGPT for data synthesis, which may pose challenges to replicability. To address this, we will open-source our generated data and code here. Future research could explore using open-source LLMs, which are increasingly powerful and comparable to proprietary models.
@@
A personalized LLM allows users to reinforce their existing beliefs and potentially amplify biased or unethical perspectives, leading to the creation of echo chambers Del Vicario et al. (2016). This can ultimately harm users by reinforcing polarized or undesirable views. To mitigate this issue, the Chain-of-Opinion (CoO) reasoning from our proposed COO involves presenting user demography or ideology group responses alongside personalized answers. Additionally, COO can encourage users to reflect on their previous viewpoints.
@@
Using ChatGPT for data synthesis.
We follow prior studies (Ray, 2023; Tan et al., 2024) to use OpenAI ChatGPT to synthesize data. Additionally, we obey OpenAI’s terms of use [^3] https://openai.com/policies/row-terms-of-use/ to use ChatGPT’s synthesized data to develop models that do not compete with OpenAI.
Misuse and responsibility for long-term societal impact.
While our method aims to align opinions with individuals, it also introduces risks of misuse, such as the propagation of harmful ideologies or manipulating human opinions. While this is not what our method is designed for, there is no way to prevent this type of misuse. We emphasize the importance of ethical alignment in deploying these systems and suggest that developers and practitioners must establish robust guidelines and oversight mechanisms to prevent misuse. Furthermore, incorporating mechanisms to monitor and audit AI-generated content following ethical norms is crucial.
We also recognize potential risks of unintended societal consequences, such as fostering group biases or undermining collective decision-making processes. We also acknowledge the potential for unintended societal consequences, such as fostering group biases or undermining collective decision-making processes. To mitigate these risks, we recommend integrating ethical safeguards, including mechanisms to monitor and audit AI use, and fostering user awareness and diverse perspectives during model training and deployment. This approach aims to minimize negative societal impacts and ensure the technology is applied constructively and equitably.
@@
Through human evaluations, we observe that our proposed method does not generate any discriminatory, insulting responses. We validate the intermediate steps of our proposed COO by human evaluation which involves manual labor. We hire annotators to score, and the hourly pay is set to \$$\$20$, which is higher than the local statutory minimum wage. Therefore, we do not anticipate any major ethical concerns raising from human evaluations.
Appendix A Implementation Details
@@
Appendix B Additional Analyses
B.1 Fine-grained Analyses
Guns Auto. vehicles Views on gender Community types Race
ChatGPT 53.87 / 57.06 45.33 / 51.25 53.21 / 59.23 43.47 / 40.88 43.06 / 43.27
ChatGPT-it 57.00 / 58.21 44.78 / 51.92 52.15 / 53.07 45.24 / 46.14 44.65 / 48.28
Mistral 44.73 / 58.12 41.72 / 53.75 40.09 / 57.78 35.45 / 42.08 41.11 / 51.44
GPT-4 60.39 / 63.37 53.22 / 50.00 63.73 / 71.43 42.86 / 47.96 55.17 / 50.57
[…]
GPT-2 large 25.34 / 30.07 26.04 / 31.96 22.66 / 24.66 23.33 / 29.00 21.67 / 22.42
FlanT5 62.08 / 60.39 55.60 / 59.61 57.45 / 62.60 45.98 / 47.20 54.23 / 61.56
FlanT5 large 60.56 / 65.45 51.52 / 58.87 60.28 / 60.64 46.21 / 49.76 49.24 / 55.82
Gender & Leadership America in 2050 Trust in science Bio. & food issues Misinformation
ChatGPT 48.27 / 52.22 46.93 / 49.46 54.93 / 56.43 52.27 / 54.75 49.33 / 60.94
ChatGPT-it 54.70 / 56.28 46.20 / 49.00 61.58 / 55.50 55.86 / 57.26 52.11 / 53.62
Mistral 50.23 / 57.87 35.14 / 47.60 51.65 / 60.37 52.78 / 58.58 50.77 / 53.85
GPT-4 65.55 / 63.03 53.71 / 45.27 61.54 / 68.46 58.03 / 61.61 52.71 / 57.26
[…]
GPT-2 large 21.50 / 28.11 27.90 / 29.27 29.61 / 30.96 28.60 / 33.52 23.34 / 30.13
FlanT5 63.98 / 66.08 55.82 / 55.00 64.22 / 63.00 61.41 / 61.56 60.49 / 60.00
FlanT5 large 58.54 / 64.00 46.54 / 51.30 61.76 / 68.43 57.34 / 63.75 51.77 / 61.97
Privacy & Surveilance Family & Relationships Economic inequality Global attitudes Political views Average
ChatGPT 53.24 / 54.29 57.22 / 61.00 45.60 / 52.43 49.60 / 45.54 56.97 / 51.15 50.22 / 52.66
ChatGPT-it 51.02 / 54.33 57.89 / 58.25 51.98 / 50.13 57.23 / 57.86 46.85 / 53.84 51.95 / 53.58
Mistral 43.31 / 58.06 47.42 / 58.50 41.87 / 51.89 42.0 / 52.76 44.13 / 53.34 44.16 / 54.40
GPT-4 47.73 / 52.27 62.50 / 63.89 63.81 / 64.76 66.67 / 63.58 62.07 / 67.82 57.98 / 59.42
[…]
GPT-2 large 29.53 / 37.30 24.19 / 28.57 22.28 / 28.21 24.87 / 32.01 23.38 / 28.46 24.94 / 30.21
FlanT5 58.36 / 58.00 64.59 / 65.60 54.06 / 57.60 63.23 / 58.04 49.08 / 52.80 55.00 / 59.62
FlanT5 large 57.92 / 61.53 64.98 / 67.02 51.66 / 58.72 57.19 / 57.95 48.55 / 56.67 54.94 / 60.13
Table 6: Fine-grained accuracy results of models with DIO-top8 Hwang et al. (2023) / Chain-of-Opinion (COO; ours).
Tab. 6 presents our fine-grained results across topics of the OpinionQA dataset Santurkar et al. (2023). Overall, COO consistently outperforms DIO-top8 across most of the OpinionQA topics. The highest absolute improvements occur in “View on gender” (+17.69 with Mistral), “Autonomous vehicles” (+13.49 with GPT-2), and “Misinformation” (+11.61 with ChatGPT). These improvements highlight the COO’s strength in helping models better handle complex, belief-based topics, particularly those involving social and political biases.
Across models, COO notably enhances performance, whether in strong models like GPT-4 or weaker ones like GPT-2. GPT-4, one of the leading LLMs for reasoning, shows marked improvement in scientific and social topics such as “Views on gender” (+7.70) and “Misinformation” (+4.55). GPT-2, usually less effective, also benefits from COO’s enhancements, with significant gains in “Views on gender” (+11.37) and “Political views” (+7.98). This further emphasizes the strength of COO in enhancing both high-performing and less capable models.
B.2 Additional Baselines
@@
We compare COO’s FEA and LLMtop-$K$ steps with two simple variants outlined in Tab. 7. Given ChatGPT and Mistral’s strong performance with just $4.79/12$ and $8.83/12$ explicit persona attributes, a crucial question arises: (1) Can comparable performance be achieved by randomly selecting $5/12$ and $9/12$ explicit persona attributes instead of relying on LLMs?. Our answer is no. The first variant, DIO-top8 + Random FEA, involves randomly selecting $5/12$ and $9/12$ explicit persona attributes. The second variant entails randomly selecting $8$ implicit persona opinions instead of using COO’s LLMtop-$K$ step. From Tab. 7, we find that randomly selecting explicit persona attributes significantly harms the performance of both models due to the removal of important attributes. Additionally, randomly selecting $8$ implicit persona opinions also adversely affects the models, particularly ChatGPT. These observations underscore the effectiveness and importance of COO’s FEA and LLMtop-$K$ steps.
B.3 Top-3 Removed Explicit Personae Attributes
Tab. 8 reveals a relatively common pattern in how three LLMs, ChatGPT, ChatGPT-it, and Mistral-7B-Instruct-v0.2, filter explicit personae across various topics. "Citizenship" and "Race" consistently emerge as the most frequently removed attributes, suggesting a deliberate effort by these models to minimize potential biases associated with these demographic factors. Additionally, ChatGPT-it’s tendency to remove "Frequency" and Mistral-7B-Instruct’s broader removal of "Education" and "Political Party" highlight model-specific strategies and comprehension in filtering personae based on topic relevance. Overall, these patterns suggest that not all explicit personae are equally relevant for predicting opinions.
@@ B.3 Top-3 Removed Explicit Personae Attributes
B.4 Consistency and Reliability of VBN versus CoT
Fig. 2 presents the consistency scores of the baseline DIO-top8 (ChatGPT) with VBN and CoT reasoning over 100 samples. Both methods show improved consistency as temperature increases, with scores rising from about 58%-82% for CoT, and 62%-88% for VBN. VBN consistently outperforms CoT across temperatures, suggesting that it is more robust and reliable compared to CoT.
B.5 Ranking Consistency for LLMtop-$K$ Step
We compute the average pairwise Overlap coefficient Vijaymeena and Kavitha (2016) $OC(A,B)=\frac{|A\cap B|}{\min(|A|,|B|)}$ across five ranking outputs generated by five input strategies Fig. 3. The performance of these strategies, evaluated on 300 random samples, is detailed in Tab. 9. Results show a negligible variance across three different random seeds, indicating that randomizing the order of implicit personae for the $LLMtop8$ step yields a relatively stable strategy.
@@ B.5 Ranking Consistency for LLMtop-$K$ Step
Model Method Seed = 2024 Seed = 5 Seed = 2000 Std
ChatGPT DIO-LLMtop8 51.03 50.95 51.11 0.0652
Mistral DIO-LLMtop8 45.86 45.55 45.36 0.2060
@@
B.6 Kendall’s Tau Scores for Ranking Agreements
@@
Figure 4: Left / Middle / Right: Ranking agreements between ChatGPT top-$K$ / ChatGPT-it / Mistral orders and semantic similarity orders. One example that has a high disagreement score is shown in Section E.2.
@@
Fig. 4 shows our ranking agreement between ChatGPT, ChatGPT-it, Mistral orders and semantic similarity orders. We observe that ChatGPT and ChatGPT-it orders have minimal monotonous relations with means approximating 0 and low standard deviations with semantic orders. More specifically, with ChatGPT, the maximum agreement is 0.6000 while the minimum is -0.5895 and the Kurtosis is -0.2173. For ChatGPT-it, the maximum is slightly lower with 0.5473, while the minimum is -0.7368 which is smaller ChatGPT, and the Kurtosis is -0.1017. Meanwhile, Mistral shows a low correlation of 0.43. These low and no correlations highlight that usefulness orders can significantly differ from the semantic similarity orders commonly used in previous studies.
B.7 Student T-test Results for Tab. 2
Model Accuracy Collapsed Accuracy
ChatGPT 4.11e-11 6.06e-13
ChatGPT-Inst. 9.97e-8 4.45e-5
GPT-4 4.23e-6 1.17e-9
Mistral 6.01e-8 4.12e-6
GPT-2-base 2.19e-69 1.82e-43
GPT-2-large 5.62e-73 6.09e-49
FlanT5-base 1.23e-19 3.19e-12
FlanT5-large 2.55e-21 1.20e-17
Table 10: The p-value computed by student t-test. We observe that all the values are significantly smaller than 0.01 verifying the significance of our improvements.
We employ the Student t-test to assess the statistical significance between COO and the best-performing baseline for each model in Tab. 2. Essentially, under the null hypothesis:
  • H0: There is no significant difference.
  • H1: There is a significant difference.
As we can see, the p-values from the tests in Tab. 10 are significantly below 0.01, indicating the significance of COO improvements.
B.8 COO’s Generalization to Other Tasks
Each step of COO aligns closely with methodologies from prior studies that have proven effective in personalized tasks. We outline them below:
  • Step 1: FEA. Filtering irrelevant user attributes to improve generation outcomes is well-established Xu et al. (2024a). For instance, Rao et al. (2011); Sakaki et al. (2014) filter gender information using classifiers, while Kim et al. (2017) focus on age, and Demszky et al. (2019) analyze political polarity. Our FEA step is thus generalizable and applicable to these tasks.
  • Step 2: LLMtop-K. Selecting the top-K most relevant historical opinions for the next prediction is conceptually related to re-ranking recommendations using LLMs Hou et al. (2024); Xu et al. (2024a) and selecting key utterances in dialogue generation Do et al. (2022). Our LLMtop-K step can likewise enhance personal chat and recommendation tasks.
  • Step 3: VBN reasoning. The Value-Belif-Norm reasoning is an adaptation of Chain-of-Thought prompting Wei et al. (2022); Kojima et al. (2022). This strategy is designed for opinion prediction but can be used for tasks where user values, beliefs, and norms are crucial features.
  • Step 4: Majority voting with a dynamic number of historical opinions. This step applies to any task where leveraging dynamic demonstrations enhances prediction accuracy.
In conclusion, COO’s steps are well-supported by prior studies and can be generalized to benefit personalized tasks.
Appendix C Prompts
C.1 Cost Analyses for API Models
Prompting costs for API models are detailed in Tab. 11. For GPT-4, COO is priced similarly to the baseline DIO-top8, while DIO-top8 + SC costs nearly twice as much. This is because we execute the FEA and LLMtop-$K$ steps of COO using ChatGPT, which is comparatively inexpensive. For ChatGPT and ChatGPT-it, COO incurs an additional $7$ to $10$ US dollars compared to DIO-top8 + SC. However, this extra cost is justified by the significant performance gains, with particularly large improvements in certain topics.
DIO-top8 DIO-top8 + CoT DIO-top8 + SC COO Model
Avg. #tokens 562.72 623.62 995.89 3227.18 ChatGPT
Total US\$ 3.01 3.73 6.82 14.05 ChatGPT
Avg. #tokens 562.72 630.58 1019.31 3418.72 ChatGPT-it
Total US\$ 3.12 3.84 7.11 20.11 ChatGPT-it
Avg. #tokens 559.27 - 1021.14* 3292.66 GPT-4
Total US\$ 91.19 - 226.15* 125.60 GPT-4
Table 11: Prompting cost analysis of COO and other baselines as of 1st Sep 2024. * denotes our estimation on $50$ samples.
C.2 Prompt Template for COO’s FEA